# Near-Field SWIPT with gMIMO in the Upper Mid-Band: Opportunities, Challenges, and the Way Forward

### v1

@@ Conclusion
The integration of near-field SWIPT with gMIMO in the upper mid-band presents a transformative pathway for achieving the energy efficiency and high data capacity required by 6G and beyond. By leveraging beamfocusing and massive spatial multiplexing, gMIMO systems can enable simultaneous energy harvesting and data transmission with unparalleled precision. These gains stem from the ability to pack more antennas into a fixed aperture at higher frequencies, enabling narrower beams and higher array gain. This allows ID users to be served more efficiently, leaving more power for energy harvesting. Additionally, near-field beamfocusing improves spatial separation, making it easier to meet ID users’ quality-of-service requirements while supporting energy delivery. This paper highlights key challenges, including the need for near-field-specific channel estimation, adaptive precoding techniques, and cost-effective array designs. Solutions such as sparse and modular arrays, as well as joint precoding of information and dedicated energy symbols, are proposed to address these challenges and enhance system performance. The case study focusing on smart factories illustrates how gMIMO arrays can be deployed in practical scenarios, offering valuable insights into optimizing SWIPT systems. Future research should focus on distributed processing, multi-modal energy harvesting, and scalable implementations to fully unlock the potential of this promising technology, paving the way for sustainable and efficient wireless communication networks.

### v2

@@ Conclusion
The integration of near-field SWIPT with gMIMO in the upper mid-band presents a transformative pathway for achieving the energy efficiency and high data capacity required by 6G and beyond. By leveraging beamfocusing and massive spatial multiplexing, gMIMO systems can enable simultaneous energy harvesting and data transmission with unparalleled precision. These gains stem from the ability to pack more antennas into a fixed aperture at higher frequencies, enabling narrower beams and higher array gain. This allows ID users to be served more efficiently, leaving more power for energy harvesting. Additionally, near-field beamfocusing improves spatial separation, making it easier to meet ID users’ quality-of-service requirements while supporting energy delivery. This paper highlights key challenges, including the need for near-field-specific channel estimation, adaptive precoding techniques, and cost-effective array designs. Solutions such as sparse and modular arrays, as well as joint precoding of information and dedicated energy symbols, are proposed to address these challenges and enhance system performance. The case study focusing on smart factories illustrates how gMIMO arrays can be deployed in practical scenarios, offering valuable insights into optimizing SWIPT systems. Future research should focus on mobile scenarios, distributed processing, multi-modal energy harvesting, and scalable implementations with experimental validation, and large-scale testbed deployment to fully unlock the potential of this promising technology, paving the way for sustainable and efficient wireless communication networks.